# Prompt Stellar and Binary Black Hole Mergers from Chemically Homogeneous Evolution in Triples

### v1

@@ 2.2.1 Assumptions
In order to evaluate the impact of a tertiary companion on the inner binary, we use analytical formulae to determine dynamical stability and assess the role of ZLK oscillations.
The dynamical instability threshold is determined from Vynatheya et al. (2022). For circular inner and outer binaries, with eccentricities $e_{\rm in}=0$ and $e_{\rm out}=0$, it simplifies to
@@
where the subscripts ‘in’ and ‘out’ refer to the inner and outer orbits, $a_{\rm{in}}$ and $a_{\rm{out}}$ are the semi-major axes, $q_{\rm out}=(m_{1}+m_{2})/m_{3}\in[10^{-2},10^{2}]$ is the mass ratio, with $m_{1}$ and $m_{2}$ being the component masses of the inner binary, $m_{3}$ being the mass of the tertiary companion, and $i_{0}$ the initial mutual inclination. In order to determine dynamical stability, we specifically consider the case where $i_{0}=0\ \rm{deg}$, as this condition results in the widest stable configurations. Our stability criteria aligns closely with others in the literature (Mardling & Aarseth, 2001; Tory et al., 2022). For our default CHE binary, we find that triples with $P_{\rm out}\lesssim 4$ d are dynamically unstable. A dynamically unstable triple system will undergo a chaotic evolution, potentially involving component exchanges, mergers, or ejections (e.g., Kiseleva et al., 1994; Iben & Tutukov, 1999; Perets & Kratter, 2012; Toonen et al., 2022). Predicting the outcome of this highly non-linear behaviour is beyond the scope of our analysis.
We assess the role of ZLK oscillations following the analytic formulae from Mangipudi et al. (2022). That study provides solutions that were tested against direct $N$-body integrations. Particularly, we want to determine, for the given triple configurations, what is the maximum eccentricity that the inner binary ($e_{\rm{max}}$) attains given an initial mutual inclination. When using this formalism, we assume the following:
@@
  • The outer eccentricity is always circular ($e_{\rm{out}}=0$). In principle, our analysis can be extended for $e_{\rm{out}}\neq 0$ if the mass ratio is unity due to a gauge-freedom of the Hamiltonian forms which keeps it axisymmetric (Tremaine, 2023; Grishin, 2024). Non-equal mass binaries with $e_{\rm{out}}\neq 0$ are susceptible to chaotic evolution and the analytical prescriptions below are not valid (e.g., Naoz et al., 2013; Stegmann et al., 2025).
  • We ignore the oblateness due to stellar rotation. The additional force rotation induces on the inner binary is subdominant to the effect of the tidal bulge in synchronous systems. Specifically, the extra precession induced from tidal bulges is $\sim 30$ times larger than from rotation-induced oblateness, even for pseudo-synchronous binaries, where there is no net angular momentum flow (Liu et al., 2015).
@@ 2.2.2 Analytic Formulae
In order to determine the maximal eccentricity ($e_{\rm max}$) of the inner binary in a triple system, we use the (implicit) analytic formulae in Eq. (17) and (24) from Mangipudi et al. (2022), which are briefly revised here.
First, we define the inner ($L_{\rm in}$) to outer ($L_{\rm out}$) angular momentum ratio
@@
where $\mu_{\rm in}=m_{1}m_{2}/(m_{1}+m_{2})$ and $\mu_{\rm out}=(m_{1}+m_{2})m_{3}/(m_{1}+m_{2}+m_{3})$ are the reduced masses of the inner and outer binaries, respectively. The value of $\eta$ is useful to quantify the hierarchy of the triple star system. In Fig. 2 we show $\eta$ while varying the tertiary mass between $1001\leq m_{3}/M_{\odot}<$ 100 and the outer orbital period between $10001<P_{\rm out}/\rm{d}<$ 1000, which defines our parameter-space region of interest. As expected, higher-mass tertiaries tend to be in the test particle limit ($\eta\ll 1)$. In contrast, low tertiary masses $m_{3}\lesssim 10\ M_{\odot}$ can result in the orbital angular momentum of the inner binary dominating over that of the outer orbit ($\eta>1$).
The analysis of triples is often simplified using the double-averaged (DA) approximation. In this approach, the gravitational effects on both the inner and outer orbits are averaged over one orbital period. Alternatively, the single-averaged (SA) approximation averages only the inner orbit over time. We define the dimensionless SA parameter
$\epsilon_{\rm SA}:=\left[\frac{a_{\rm{in}}}{a_{\rm{out}}(1-e_{\rm{out}}^{2})}\right]^{3/2}\left[\frac{m_{3}^{2}}{(m_{1}+m_{2})(m_{1}+m_{2}+m_{3})}\right]^{1/2}$ (3)
@@
We proceed to assess the relative importance of short-range forces, which can also be encapsulated with dimensionless parameters. We follow the definitions of Liu et al. (2015) and write the relative strength of the apsidal precession due to general relativity (GR) and tides as
@@
respectively, where $r_{g}=G(m_{1}+m_{2})/c^{2}$ is the combined Schwarzschild radius and $R_{c}$ is the convective radius of the star, which is generally smaller than the stellar surface ($R$). These parameters, $\epsilon_{\rm GR}$ and $\epsilon_{\rm Tide}$, quantify the relative importance of the short-range potential terms with respect to the quadruple potential leading to the ZLK effect (Liu et al., 2015). For small $\epsilon_{\rm GR}\ll 1$ and $\epsilon_{\rm Tide}\ll 1$ the orbital evolution is approximated by Newtonian point masses and the eccentricity determined exclusively by the initial mutual inclination. For large $\epsilon_{\rm GR}\gg 1$ and $\epsilon_{\rm Tide}\gg 1$, additional parameters become relevant, and the eccentricity evolution may be suppressed. Consequently, the binary system may evolve as if isolated, unaffected by the tertiary component.
In Fig. 3, we show these dimensionless parameters for our default CHE binary (§2) at the ZAMS. In the top and bottom panels we display $\epsilon_{\rm GR}$ and $\epsilon_{\rm Tide}$, respectively. The dynamic range of these two parameters is broad, spanning several orders of magnitude across our triple orbital parameter space of interest. We find that, among short-range forces, tidal effects dominate the evolution at the ZAMS.
@@
where $g_{\rm max}=\cos i_{0}+\eta e_{\rm max}^{2}/2$, and $j_{\rm min}=\sqrt{1-e_{\rm max}^{2}}$. Eq. (6) can be solved for $e_{\rm max}$ in terms of $\epsilon_{\rm SA},\eta,i_{0}$, $\epsilon_{\rm GR}$ and $\epsilon_{\rm{tide}}$, which are determined from initial conditions.
@@
We note that Mangipudi et al. (2022) considers highly eccentric regimes. In our analysis, we will find that the latter term for $\delta e$ is not significant during the stellar lifetime. However, it may become relevant after the formation of a BBH, once tidal effects are no longer present.
@@ 2.2.3 Merger Criteria
The ZLK mechanism results in a periodic exchange between the mutual inclination and the eccentricity of the inner binary. However, due to conservation of angular momentum, the semi-major axis of the inner binary remains constant. While the semi-major axis remains constant, changes in eccentricity affect the pericenter distance $r_{\rm p}=a(1-e)$.
For a stellar binary, the pericenter is crucial in order to determine whether or not a binary remains detached or if at least one components fills their Roche lobe. In order for Roche-lobe overflow to begin, the radius of the star must be larger than the Roche radius ($R_{\rm RL}$). Some binary systems, sometimes referred to as contact binaries, can have both stars filling their Roche lobes while remaining in a dynamically stable configuration (Almeida et al., 2015). CHE binaries are predicted to exist in such contact configurations (Marchant et al., 2016).
We follow the criteria from Marchant et al. (2016) in order to determine whether or not a contact binary will lead to a merger. In this criteria, if both stars overflow their outer Lagrangian point ($L_{2}$), the mass lost through this point will carry high specific angular momentum, leading to a stellar merger. The criterion $R>L_{2}$ can be rewritten in terms of the Roche radius using $L_{2}\approx 1.32R_{\rm RL}$. We calculate the Roche radius at pericenter as $R_{\rm RL}=r_{\rm p}f(q)$, where $f(q)$ is a function dependant on the mass ratio. We utilize the fitting formula from Eggleton (1983), based on a circular orbit, to determine $f(q)$. For a mass ratio of unity, like in our default CHE binary, the criterion is approximately simplified to
@@
and a merger occurs if the stars make contact at pericenter.
Figure 4: Characteristic von Zeipel-Lidov-Kozai timescale ($\tau_{\rm ZLK}$) for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity $Z=0.00042$, evaluated at the zero-age main sequence. The lines indicate different relevant timescales for the inner binary. The white dotted, black dashed and black solid lines indicate the circularisation ($\tau_{\rm circ}$), synchronisation ($\tau_{\rm sync}$), and apsidal tidal ($\tau_{\rm aps}$) timescales, respectively. Dark grey indicates dynamically unstable triples.
@@ 2.3.1 ZLK Timescale
To estimate the timescale over which the ZLK mechanism can influence the eccentricity of the inner binary, we follow Liu et al. (2015) and define the characteristic ZLK timescale [^1] Eq. (10) is an order of magnitude estimate of the secular timescale, which is used to measure the different strength of the terms in the total Hamiltonian (Liu et al., 2015). The more accurate estimate requires integrating the action $G_{\rm in}\propto\sqrt{1-e_{\rm in}^{2}}$ over a closed orbit in phase space (Antognini, 2015; Grishin, 2024), which differs by an order unity from Eq. (10). as
@@
In Fig, 4, we show $\tau_{\rm ZLK}$ in our parameter-space region of interest. The largest values for the ZLK timescale are $\tau_{\rm ZLK}\sim 10^{4}$ yr, which is significantly shorter than the lifetime of the component stars of the binary ($\approx 5$ Myr, see Fig. 1). We proceed to estimate the relevant tidal timescales, which are solely determined by the structure of the inner binary, in order to compare them with the ZLK timescale, which depends on the triple orbit.
@@ 2.3.2 Apsidal Tidal Timescale
For our default CHE binary at the ZAMS (§2.1.1), this results in an apsidal tidal timescale of $\tau_{\rm{aps}}\approx 0.1\ \rm yr$. By definition, the region where $\tau_{\rm ZLK}\approx\tau_{\rm aps}$ corresponds to $\epsilon_{\rm Tide}\approx 1$, which delimits the region where the static tide becomes strong enough to suppress the ZLK mechanism ($\epsilon_{\rm Tide}\gtrsim 1$, Fig. 3).
@@ 2.3.3 Tidal Dissipation Timescale
In a stellar binary system, tidal dissipation eventually aligns both stellar spins with the orbit, synchronizes the components with the orbital motion, and makes the orbit circular (e.g., Zahn, 2008, and references therein). The efficiency of tidal dissipation, however, is highly dependent on the specific dissipation mechanism (Zahn, 1977) and is also very sensitive to the assumptions made about the stellar structure (e.g., Siess et al., 2013; Qin et al., 2018). In binary star systems with stars that have radiative envelopes, such as those in our CHE binary, the primary dissipation mechanism is through the dynamical tide (Zahn, 1975, 1977).
We follow Hurley et al. (2002) and define the circularization and synchronization timescales via the dynamical tide for a near-circular binary as
@@
following the models from Claret & Gimenez (1989).
For our default CHE binary at the ZAMS (§2.1.1), this results in a circularisation timescale of $\tau_{\rm{circ}}\approx 1500\ \rm yr$ and a synchronisation timescale of $\tau_{\rm{sync}}\approx 23\ \rm yr$.
@@ 2.3.4 Comparison
In Fig. 4 we can see how the ZLK timescale compares to the apsidal tidal timescale (solid black line) and the tidal dissipation timescales, specifically the circularization (dotted white line) and synchronisation (dashed black line) timescales. We differentiate between four main regimes:
@@
We proceed to explicitly show the impact of short-range forces and the ZLK mechanism on the orbital configuration of the inner binary.
@@ 3 Results
Figure 5: Maximal inner eccentricity ($e_{\rm max}$) dynamically induced, via the von Zeipel-Lidov-Kozai (ZLK) mechanism, by a tertiary for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity $Z=0.00042$, evaluated at the zero-age main sequence. We use different colours to highlight various regions of interest. Dark grey indicates dynamically unstable triples. If the tertiary is unable to torque the inner binary, the latter remains circular ($e_{\rm max}\approx 0$); white regions indicate this regime. We explore the impact of different assumptions (§2.2.2). The left and right columns show the results using the double-averaged (DA) and single-averaged (SA) approximations, respectively. From top to bottom, the models incorporate additional terms that account for relativistic apsidal precession (GR), apsidal precession from the tidal equilibrium bulge in the inner binary (Tides), and both.
For our default CHE binary (§2), we run a grid in $m_{3}$, $a_{\rm out}$, and $i_{0}$. For $m_{3}$ and $a_{\rm out}$, we create a 250$\times$250 $\log$-uniform grid in the range $1001\leq m_{3}/M_{\odot}<$ 100 and $2a_{\rm in}\leq a_{\rm out}<200a_{\rm in}$, which covers the parameter-space of interest between $10001<P_{\rm out}/\rm{d}<$ 1000. For $i_{0}$, we create a uniform grid with 181 gridpoints between $\cos(i_{0})=\{-1,1\}$, making sure to include $\cos(i_{0})=0$. We numerically solve Eq. (6) for $e_{\rm max}$ using the bisection method, which then allows us to calculate the fraction of stellar mergers (see §2.2.3 for our criteria).
Here, we present the results of our analysis focusing on the following aspects: the maximum eccentricity under various combinations of short-range forces (§3.1), the predicted fraction of stellar mergers (§3.2), and the broader implications for populations of stellar binaries and BBHs (§3.3).

### v2

@@ 2.2.1 Assumptions
To evaluate the impact of a tertiary companion on the inner binary, we used analytical formulae to determine dynamical stability and assess the role of ZLK oscillations.
The dynamical instability threshold is determined from Vynatheya et al. (2022). For circular inner and outer binaries, with eccentricities $e_{\rm in}=0$ and $e_{\rm out}=0$, it can be simplified to
@@
where the subscripts ‘in’ and ‘out’ refer to the inner and outer orbits, $a_{\rm{in}}$ and $a_{\rm{out}}$ are the semi-major axes, $q_{\rm out}=(m_{1}+m_{2})/m_{3}\in[10^{-2},10^{2}]$ is the mass ratio, with $m_{1}$ and $m_{2}$ being the component masses of the inner binary and $m_{3}$ being the mass of the tertiary companion, and $i_{0}$ the initial mutual inclination. In order to determine dynamical stability, we specifically considered the case where $i_{0}=0\ \rm{deg}$, as this condition results in the widest stable configurations. Our stability criteria is closely aligned with other works in the literature (Mardling & Aarseth, 2001; Tory et al., 2022). For our default CHE binary, we find that triples with $P_{\rm out}\lesssim 4$ d are dynamically unstable. A dynamically unstable triple system will undergo a chaotic evolution, potentially involving component exchanges, mergers, or ejections (e.g. Kiseleva et al., 1994; Iben & Tutukov, 1999; Perets & Kratter, 2012; Toonen et al., 2022). Predicting the outcome of this highly non-linear behaviour is beyond the scope of our analysis.
We assessed the role of ZLK oscillations following the analytic formulae from Mangipudi et al. (2022). That study provides solutions that were tested against direct $N$-body integrations. In particular, we sought to determine, for the given triple configurations, the maximum eccentricity that the inner binary ($e_{\rm{max}}$) can attain given an initial mutual inclination. When using this formalism, we assume the following:
@@
  • The outer eccentricity is always circular ($e_{\rm{out}}=0$). In principle, our analysis can be extended for $e_{\rm{out}}\neq 0$ if the mass ratio is unity due to a gauge-freedom of the Hamiltonian forms, which keeps it axisymmetric (Tremaine, 2023; Grishin, 2024). Non-equal mass binaries with $e_{\rm{out}}\neq 0$ are susceptible to chaotic evolution and the analytical prescriptions below are not valid (e.g. Naoz et al., 2013; Stegmann et al., 2025).
  • We ignored the oblateness due to stellar rotation. The additional force rotation induces on the inner binary is subdominant to the effect of the tidal bulge in synchronous systems. Specifically, the extra precession induced from tidal bulges is $\sim 30$ times larger than from rotation-induced oblateness, even for pseudo-synchronous binaries, where there is no net angular momentum flow (Liu et al., 2015).
@@ 2.2.2 Analytic formulae
To determine the maximal eccentricity ($e_{\rm max}$) of the inner binary in a triple system, we used the (implicit) analytic formulae in Eq. (17) and (24) from Mangipudi et al. (2022), which are briefly revised here. First, we define the inner ($L_{\rm in}$) to outer ($L_{\rm out}$) angular momentum ratio,
@@
where $\mu_{\rm in}=m_{1}m_{2}/(m_{1}+m_{2})$ and $\mu_{\rm out}=(m_{1}+m_{2})m_{3}/(m_{1}+m_{2}+m_{3})$ are the reduced masses of the inner and outer binaries, respectively. The value of $\eta$ is useful to quantify the hierarchy of the triple star system. In Fig. 2, we show $\eta$ while varying the tertiary mass between $1001\leq m_{3}/M_{\odot}<$ 100 and the outer orbital period between $10001<P_{\rm out}/\rm{d}<$ 1000, which defines our parameter-space region of interest. As expected, higher-mass tertiaries tend to be in the test particle limit ($\eta\ll 1)$. In contrast, low tertiary masses $m_{3}\lesssim 10\ M_{\odot}$ can result in the orbital angular momentum of the inner binary dominating over that of the outer orbit ($\eta>1$).
The analysis of triples is often simplified using the double-averaged (DA) approximation. In this approach, the gravitational effects on both the inner and outer orbits are averaged over one orbital period. Alternatively, the single-averaged (SA) approximation averages only the inner orbit over time. We define the dimensionless SA parameter as
$\epsilon_{\rm SA}:=\left[\frac{a_{\rm{in}}}{a_{\rm{out}}(1-e_{\rm{out}}^{2})}\right]^{3/2}\left[\frac{m_{3}^{2}}{(m_{1}+m_{2})(m_{1}+m_{2}+m_{3})}\right]^{1/2},$ (3)
@@
We proceeded to assess the relative importance of short-range forces, which can also be encapsulated with dimensionless parameters. We followed the definitions of Liu et al. (2015) and write the relative strength of the apsidal precession due to general relativity (GR) and tides as
@@
respectively, where $r_{g}=G(m_{1}+m_{2})/c^{2}$ is the combined Schwarzschild radius and $R_{c}$ is the convective radius of the star, which is generally smaller than the stellar surface ($R$). These parameters, $\epsilon_{\rm GR}$ and $\epsilon_{\rm Tide}$, quantify the relative importance of the short-range potential terms with respect to the quadruple potential leading to the ZLK effect (Liu et al., 2015). For small $\epsilon_{\rm GR}\ll 1$ and $\epsilon_{\rm Tide}\ll 1$ the orbital evolution is approximated by Newtonian point masses and the eccentricity determined exclusively by the initial mutual inclination. For large values $\epsilon_{\rm GR}\gg 1$ and $\epsilon_{\rm Tide}\gg 1$, additional parameters become relevant and the eccentricity evolution may be suppressed. Consequently, the binary system may evolve as if isolated, unaffected by the tertiary component.
In Fig. 3 we show these dimensionless parameters for our default CHE binary (Sect. 2) at the ZAMS. In the top and bottom panels we display $\epsilon_{\rm GR}$ and $\epsilon_{\rm Tide}$, respectively. The dynamic range of these two parameters is broad, spanning several orders of magnitude across our triple orbital parameter space of interest. We find that, among short-range forces, tidal effects dominate the evolution at the ZAMS.
@@
where $g_{\rm max}=\cos i_{0}+\eta e_{\rm max}^{2}/2$, and $j_{\rm min}=\sqrt{1-e_{\rm max}^{2}}$. Eq. (6) can be solved for $e_{\rm max}$ in terms of $\epsilon_{\rm SA},\eta,i_{0}$, $\epsilon_{\rm GR}$, and $\epsilon_{\rm{tide}}$, which are determined from initial conditions.
@@
We note that Mangipudi et al. (2022) considers highly eccentric regimes. In our analysis, we ultimately found that the latter term for $\delta e$ is not significant during the stellar lifetime. However, it may become relevant after the formation of a BBH, once tidal effects are no longer present.
@@ 2.2.3 Merger criteria
The ZLK mechanism results in a periodic exchange between the mutual inclination and the eccentricity of the inner binary. However, due to conservation of angular momentum, the semi-major axis of the inner binary remains constant. While the semi-major axis remains constant, changes in eccentricity affect the pericentre distance $r_{\rm p}=a(1-e)$.
For a stellar binary, the pericentre is crucial in order to determine whether or not a binary remains detached or if at least one components fills their Roche lobe. For Roche-lobe overflow to begin, the radius of the star must be larger than the Roche radius ($R_{\rm RL}$). Some binary systems, sometimes referred to as contact binaries, can have both stars filling their Roche lobes while remaining in a dynamically stable configuration (Almeida et al., 2015). CHE binaries are predicted to exist in such contact configurations (Marchant et al., 2016).
We follow the criteria from Marchant et al. (2016) in order to determine whether or not a contact binary will lead to a merger. In this criteria, if both stars overflow their outer Lagrangian point ($L_{2}$), the mass lost through this point will carry high specific angular momentum, leading to a stellar merger. The criterion $R>L_{2}$ can be rewritten in terms of the Roche radius using $L_{2}\approx 1.32R_{\rm RL}$. We calculate the Roche radius at pericentre as $R_{\rm RL}=r_{\rm p}f(q)$, where $f(q)$ is a function that is dependent on the mass ratio. We utilised the fitting formula from Eggleton (1983), based on a circular orbit, to determine $f(q)$. For a mass ratio of unity, as in our default CHE binary, the criterion is approximately simplified to
@@
and a merger occurs if the stars make contact at pericentre.
Figure 4: Characteristic ZLK timescale ($\tau_{\rm ZLK}$) for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity of $Z=0.00042$, evaluated at the ZAMS. The lines indicate different relevant timescales for the inner binary. The dotted white, dashed black, and solid black lines indicate the circularisation ($\tau_{\rm circ}$), synchronisation ($\tau_{\rm sync}$), and apsidal tidal ($\tau_{\rm aps}$) timescales, respectively. Dark grey indicates dynamically unstable triples.
@@ 2.3.1 ZLK timescale
To estimate the timescale over which the ZLK mechanism can influence the eccentricity of the inner binary, we follow Liu et al. (2015) and define the characteristic ZLK timescale [^1] Equation (10) is an order-of-magnitude estimate of the secular timescale, which is used to measure the different strength of the terms in the total Hamiltonian (Liu et al., 2015). The more accurate estimate requires integrating the action $G_{\rm in}\propto\sqrt{1-e_{\rm in}^{2}}$ over a closed orbit in phase space (Antognini, 2015; Grishin, 2024), which differs by an order unity from Eq. (10). as
@@
In Fig, 4 we show $\tau_{\rm ZLK}$ in our parameter-space region of interest. The largest values for the ZLK timescale are $\tau_{\rm ZLK}\sim 10^{4}$ yr, which is significantly shorter than the lifetime of the component stars of the binary ($\approx 5$ Myr; see Fig. 1). We proceed to estimate the relevant tidal timescales, which are solely determined by the structure of the inner binary, in order to compare them with the ZLK timescale, which depends on the triple orbit.
@@ 2.3.2 Apsidal tidal timescale
For our default CHE binary at the ZAMS (Sect. 2.1.1), this results in an apsidal tidal timescale of $\tau_{\rm{aps}}\approx 0.1\ \rm yr$. By definition, the region where $\tau_{\rm ZLK}\approx\tau_{\rm aps}$ corresponds to $\epsilon_{\rm Tide}\approx 1$, which delimits the region where the static tide becomes strong enough to suppress the ZLK mechanism ($\epsilon_{\rm Tide}\gtrsim 1$, Fig. 3).
@@ 2.3.3 Tidal dissipation timescale
In a stellar binary system, tidal dissipation eventually aligns both stellar spins with the orbit, synchronises the components with the orbital motion, and makes the orbit circular (e.g. Zahn, 2008, and references therein). The efficiency of tidal dissipation, however, is highly dependent on the specific dissipation mechanism (Zahn, 1977) and is also very sensitive to the assumptions made about the stellar structure (e.g. Siess et al., 2013; Qin et al., 2018). In binary star systems with stars that have radiative envelopes, such as those in our CHE binary, the primary dissipation mechanism is through the dynamical tide (Zahn, 1975, 1977).
We follow Hurley et al. (2002) and define the circularisation and synchronisation timescales via the dynamical tide for a near-circular binary as
@@
following the models from Claret & Gimenez (1989). For our default CHE binary at the ZAMS (Sect. 2.1.1), this results in a circularisation timescale of $\tau_{\rm{circ}}\approx 1500\ \rm yr$ and a synchronisation timescale of $\tau_{\rm{sync}}\approx 23\ \rm yr$.
@@ 2.3.4 Comparison
In Fig. 4 we can see how the ZLK timescale compares to the apsidal tidal timescale (solid black line) and the tidal dissipation timescales specifically, the circularisation (dotted white line) and synchronisation (dashed black line) timescales. We differentiate between four main regimes:
@@
We explicitly demonstrate the impact of short-range forces and the ZLK mechanism on the orbital configuration of the inner binary in the next section.
@@ 3 Results
Figure 5: Maximal inner eccentricity ($e_{\rm max}$) dynamically induced, via the ZLK mechanism, by a tertiary for a triple system with an arbitrary tertiary mass ($m_{3}$) and outer orbital period ($P_{\rm out}$). The system includes an inner circular binary with masses $m_{1}=m_{2}=55\ M_{\odot}$ and an orbital period of 1.1 d, at a metallicity $Z=0.00042$, evaluated at the ZAMS. We use different colours to highlight various regions of interest. Dark grey indicates dynamically unstable triples. If the tertiary is unable to torque the inner binary, the latter remains circular ($e_{\rm max}\approx 0$); white regions indicate this regime. We explore the impact of different assumptions (Sect. 2.2.2). The left and right columns show the results using the double-averaged (DA) and single-averaged (SA) approximations, respectively. From top to bottom, the models incorporate additional terms that account for relativistic apsidal precession (GR), apsidal precession from the tidal equilibrium bulge in the inner binary (tides), and both.
For our default CHE binary (Sect. 2), we ran a grid in $m_{3}$, $a_{\rm out}$, and $i_{0}$. For $m_{3}$ and $a_{\rm out}$, we created a 250$\times$250 $\log$-uniform grid in the range $1001\leq m_{3}/M_{\odot}<$ 100 and $2a_{\rm in}\leq a_{\rm out}<200a_{\rm in}$, which covers the parameter-space of interest between $10001<P_{\rm out}/\rm{d}<$ 1000. For $i_{0}$, we created a uniform grid with 181 grid points between $\cos(i_{0})=\{-1,1\}$, making sure to include $\cos(i_{0})=0$. We numerically solved Eq. (6) for $e_{\rm max}$ using the bisection method, which then allowed us to calculate the fraction of stellar mergers (see Sect. 2.2.3 for our criteria).
Here, we present the results of our analysis focusing on the following aspects: the maximum eccentricity under various combinations of short-range forces (Sect. 3.1), the predicted fraction of stellar mergers (Sect. 3.2), and the broader implications for populations of stellar binaries and BBHs (Sect. 3.3).